# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v2

@@ 4.2.2. Node Alignment
The Aligner module utilizes the summarized reasoning chain to identify and align relevant nodes within the graph, effectively pruning redundant nodes and edges by filtering out unrelated information at the node level. Specifically, we employ a GNN (e.g., GraphTransformer (Shi et al., 2020) or GAT (Veličković et al., 2018)), to encode the structural information of the graph. The GNN produces node-level embeddings $\boldsymbol{n}_{g}$ based on the input graph:
(8) $\boldsymbol{n}_{g}=\text{GNN}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d},$
@@
where $|\mathcal{V}|$ is the number of node and $d$ is feature dim. For the reasoning text, we employ SBERT to encode the textual description into a representation $\boldsymbol{r}_{s}$ that captures its semantic meaning:
(9) $\boldsymbol{r}_{s}=\text{SBERT}(t_{\text{reasoning}})\in\mathbb{R}^{d_{s}}$
@@
We concatenate the embedding of each node in the subgraph with the query embedding. The concatenated embeddings are then passed through an MLP module, which generates a predicted importance score for each node. Finally, these scores are transformed into probability distributions using the softmax function, producing importance scores for both prediction and reasoning.
(10) $\boldsymbol{p}_{\text{prediction}}$ $=\text{Softmax}(\text{MLP}([\boldsymbol{n}_{g},\boldsymbol{q}^{\text{expand}}]))\in\mathbb{R}^{|\mathcal{V}|},$
(11) $\boldsymbol{p}_{\text{reasoning}}$ $=\text{Softmax}(\text{sim}(\boldsymbol{r}_{s},\boldsymbol{n}))\in\mathbb{R}^{|\mathcal{V}|}$
@@
where $[,]$ means concat operation, $\boldsymbol{q}^{\text{expand}}$ is the query embedding broadcasted across all nodes. To align the node distribution, we minimize the Kullback–Leibler (KL) divergence (Kullback and Leibler, 1951) between predicted probabilities $\boldsymbol{p}_{\text{prediction}}$ and the probabilities $\boldsymbol{p}_{\text{reasoning}}$. The KL divergence loss for a subgraph is given by:
(12) $\mathcal{L}_{\text{NA}}=\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\boldsymbol{p}_{\text{reasoning}}(i)\log\frac{\boldsymbol{p}_{\text{reasoning}}(i)}{\boldsymbol{p}_{\text{prediction}}(i)}$
@@
4.2.3. Representation Alignment
To bridge the representation gap between graph structures and language-based descriptions, by treating the text representation derived from reasoning chain as the target label, Aligner module aligns the graph and text embeddings to encourage semantic consistency. We apply a mean pooling operation across the node embeddings $\boldsymbol{n}_{g}$ to obtain a unified graph-level representation $\boldsymbol{r}_{g}$:
(13) $\boldsymbol{r}_{g}=\text{POOL}(\boldsymbol{n}_{g})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\boldsymbol{n}_{g}(v)\in\mathbb{R}^{d}$
@@ 4.2.3. Representation Alignment
(14) $\boldsymbol{\hat{r}}_{s}$ $=\text{MLP}(\boldsymbol{r}_{s})\in\mathbb{R}^{d_{t}},$
(15) $\boldsymbol{\hat{r}}_{g}$ $=\text{MLP}(\boldsymbol{r}_{g})\in\mathbb{R}^{d_{t}}$
@@
The contrastive loss for representation alignment from $\boldsymbol{\hat{r}}_{g}$ to $\boldsymbol{\hat{r}}_{s}$ is defined as:
(16) $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s})=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{j})/\tau\right)}\right]$
@@
where, $N$ is the batch size, $(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})$ is the $i$-th positive (graph-text) pair in the batch, $\tau$ is a temperature parameter to control the sharpness. Similarly, we can obtains the loss $\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})$ from $\boldsymbol{\hat{r}}_{s}$ to $\boldsymbol{\hat{r}}_{g}$. The final representation alignment loss is obtained by:
(17) $\mathcal{L}_{RA}=\frac{1}{2}(\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})+\mathcal{L}_{RA}(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s}))$
To achieve an optimized Graph Aligner, we perform joint optimization of node and representation Alignment. The total loss for the Graph Aligner is defined as:
(18) $\mathcal{L}_{\text{Aligner}}=\mathcal{L}_{\text{RA}}+\mathcal{L}_{\text{NA}}$
@@
The parameters of GNN encoder are jointly optimized using the loss $\mathcal{L}_{\text{Aligner}}$ with a specified training step, which reflects degree of alignment. We evaluate the impact of the alignment degree in Figure 4 and the effectiveness in Section 5.7.
4.2.4. Graph Pruning and Representation Generation
We use the trained Graph Aligner to prune irrelevant nodes, and produce graph representations that are aligned with language. The Aligner performs graph pruning across the entire dataset. To start, we introduce a hyperparameter: the number of seed nodes, denoted as $n_{\text{seed}}$. For the graph’s node-level output, $\boldsymbol{n}_{\text{prediction}}$, the top $n_{\text{seed}}$ nodes are selected as seed nodes. Using the first-order neighbors of these seed nodes, we expand the adjacent nodes and edges. This process yields the newly constructed subgraph, denoted as $\mathcal{G}_{\text{Aligner}}=(\mathcal{V}_{\text{Aligner}},\mathcal{E}_{\text{Aligner}})$. What is more, we also utilize the trained Aligner across the entire dataset to produce the graph representation $\boldsymbol{r}_{g}$. The aligned subgraph $\mathcal{G}_{\text{Aligner}}$ serves as the input to LLM, enabling more efficient and accurate generation results. The efficiency of pruning is evaluated in section 5.8.
4.3. Graph Generator
The Aligner module enables efficient representation alignment and pruning of irrelevant knowledge, greatly facilitating integration with LLMs. To generate the final answer, we concatenate the text of graph $\mathcal{G}_{\text{Aligner}}$ with query tokens and input them into the LLM’s TextEmbedder ((The token embedding of LLM itself). This yields text tokens embedding: $\boldsymbol{r}_{t}=\text{TextEmbedder}([t_{q},\mathcal{G}_{\text{Aligner}}])\in\mathbb{R}^{(T_{q}+T_{g})\times d_{t}}$, where $T_{q}$ and $T_{g}$ mean the token length of query and graph. $\boldsymbol{r}_{t}$ is then concatenated with the previously learned graph embedding $\boldsymbol{r}_{g}$. The conditional probability of generating answer $Y$ given the aligned graph $\mathcal{G}_{\text{Aligner}}$ and the query $t_{q}$ is defined as:
@@ 4.3. Graph Generator
(19) $p_{\theta}(Y|t_{q},\mathcal{G}_{\text{Aligner}})=\prod_{i=1}^{m}p_{\theta}(y_{i}|[\text{MLP}(\boldsymbol{r}_{g}),\boldsymbol{r}_{t}],y_{<i})$
@@
To enhance efficiency of training and deployment, we fine-tune generator using parameter-efficient tuning method, like LoRA (Hu et al., 2021).
5. Experimental Settings
Table 1. Statistics of GraphQA Benchmark (He et al., 2024).
Dataset #Training #Validation #Test Task Type
ExplaGraphs 1,659 553 554 Commonsense reasoning
SceneGraphs 59,978 19,997 20,025 Visual scene understanding
WebQSP 2,826 245 1,628 Knowledge-based QA
5.1. Datasets and Metrics.
Following G-Retriever (He et al., 2024), we conducted experiments on the GraphQA benchmark (He et al., 2024), which includes ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). The dataset statistics are summarized in Table 1, with a data split of train:validation:test = 6:2:2.
5.2. Baselines.
5.2.1. Inference-only.
  • Zero-CoT (Wei et al., 2022): enhances zero-shot reasoning by appending the phrase “Let’s think step by step.”, which encourages the model to explicitly generate intermediate reasoning steps.
  • CoT-BAG (Wang et al., 2024): adds “Let’s construct a graph with the nodes and edges first.” after providing the textual graph description,
  • KAPING (Baek et al., 2023): augments LLMs with knowledge graph facts at the prompt level, enabling more accurate QA without requiring model fine-tuning retrieving relevant graph triples.
  • ToG (Sun et al., [n. d.]): integrates LLMs with KGs to support deep and reliable reasoning through interactive exploration of entities and relations.
5.2.2. Raw Fine-tuning.
In this setting, following (He et al., 2024), we fine-tune the LLM via parameter-efficient methods, without applying advanced reranking techniques. Two widely used approaches are included:
  • Prompt Tuning (Lester et al., 2021): learns a set of continuous prompt embeddings that guide the frozen LLM during inference, enabling lightweight adaptation to specific tasks.
  • LoRA (Hu et al., 2021): introduces low-rank adaptation layers into transformer weights, achieving efficient fine-tuning with significantly reduced computational and memory costs.
5.2.3. Reranker-based.
  • GTE (General Textual Embedding) (Li et al., 2023b): developed by Alibaba DAMO Academy, with two variants (gte-base, 109M parameters; gte-large, 335M parameters), trained on large-scale relevance pairs to improve retrieval accuracy,
  • BGE (BAAI General Embedding) (Xiao et al., 2024): cross-encoder models with two variants (bge-reranker-base, 278M parameters; bge-reranker-large, 560M parameters), optimized for retrieval-augmented generation, offering high accuracy at the cost of efficiency,
  • G-RAG and G-RAG-RL (Dong et al., 2024): reranker models that leverage GNNs to incorporate document connections and semantic cues from abstract meaning representation graphs for context-aware ranking.
5.2.4. GNN-based.
  • GraphToken (Perozzi et al., 2024): encodes graph structures as explicit prompts for LLMs, improving performance on graph reasoning tasks,
  • GNP (Tian et al., 2024): a plug-and-play framework that applies a GNN encoder with cross-modality pooling to enrich LLMs with knowledge graph information,
  • G-Retriever (He et al., 2024): introduces the GraphQA benchmark and employs retrieval-augmented generation with soft prompting, improving graph-based question answering, comprehension, and reducing hallucinations,
  • GRAG (Hu et al., 2024): proposes a divide-and-conquer strategy for efficient textual subgraph retrieval and combines text and graph views within LLMs for graph context-aware generation.
5.3. Implementation Details.
We use GraphTransformer (Shi et al., 2020), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2016) as GNN encoders, and Llama-2-7b-hf, Llama-2-13b-hf and Llama-3.1-8b (Touvron et al., 2023) as generators. For the retrieval process, we use the same retrieval results obtained through the PCST algorithm across all baselines to ensure a fair comparison. Both reranker-based and GNN-based methods apply LoRA for fine-tuning, except for GraphToken, whose results are from G-Retriever (He et al., 2024). All methods are compared using the same training hyperparameters where applicable, ensuring fair comparability (e.g., GNN layers and LoRA rank). In the Aligner module, we explore two hyperparameters: alignment degree and the number of seed nodes ($n_{\text{seed}}$), with the analysis shown in Figure 7. To summarize the reasoning chain, we employ Llama-3.1-70B-Instruct (Dubey et al., 2024). The models were trained on the training set, with optimal hyperparameters and early stopping determined using the validation set. Results are reported on the test set. All experiments were run on two NVIDIA A100 GPUs (80GB each).
To implement baselines, for GNP (Tian et al., 2024), we implemented the Graph Neural Prompting module within our framework, including components such as the GNN encoder, cross-modality pooling, and domain projector. For G-RAG and G-RAG-RL (Dong et al., 2024), we adopted their ranking approach, combining cross-entropy loss with pairwise ranking loss. However, since treating documents as nodes was infeasible in our case, we instead treated entities as nodes and employed the same GNN encoder as our method. We implemented GTE (Li et al., 2023b) and BGE-reranker (Xiao et al., 2024) using their official open-source models. Nodes and triples are ranked by query similarity and, as in our method, fed into the LLM for generation. For the $\text{GRAG}_{PCST}$ model (Hu et al., 2024), we reproduced its experiments. However, to ensure a fair evaluation, we standardized the retrieval process by using the PCST method for retrieval. This allowed us to directly compare it with their graph encoder approach.
5.4. Main Results
Table 2. Performance comparison using the same retrieval settings for all methods. Llama-2-7b-hf is used as generator. The table reports the mean and standard deviation (std) results across three random seeds. For methods marked with $\dagger$’, we reproduce the results on WebQSP to report both F1 and Accuracy. Results for methods marked with $\ddagger$’ are taken directly from (He et al., 2024). The best results are highlighted in bold, and the second-best results are underlined. ‘Improvement $(\triangle)$’ represents the gain of Align-GRAG over the second-best baseline. ‘*’ indicates the statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over the best baseline. $\uparrow$: higher is better.
@@ 5.4. Main Results
In this section, we conduct extensive experiments on the ExplaGraphs, SceneGraphs, and WebQSP datasets, comparing the performance of our method with 16 baseline methods. All baselines are implemented using the same retrieved text, and the same backbones, Llama2-7b (Touvron et al., 2023) and GraphTransformer (Shi et al., 2020). As shown in Table 2, the overall results demonstrate the effectiveness of Align-GRAG, which achieves performance improvements across all metrics on three datasets. Notably, on WebQSP dataset, Align-GRAG achieves a remarkable improvement in Accuracy compared to the second-best method. This improvement can be attributed to the innovative design of Align-GRAG, particularly its alignment-based pruning and optimization strategies for graph structures, showcasing its potential in complex graph reasoning and knowledge-based question answering tasks.
Comparing with Inference-only methods. Inference-only methods rely solely on the reasoning capabilities of LLMs without incorporating task-specific optimization, which often leads to subpar performance. Specifically, the Zero-shot approach, which does not depend on any carefully constructed prompts, suffers from limited performance due to its inability to leverage task-specific knowledge. In contrast, ToG achieves superior performance by iteratively executing beam search on KG, enabling the discovery of promising reasoning paths. These results underscore ToG’s effectiveness within inference-only approaches.
Comparing with raw fine-tuning methods. LoRA significantly outperforms Prompt Tuning and also surpasses inference-only methods. This highlights the effectiveness of task-specific optimization in graph QA scenarios. However, it still lags behind reranker-based and GNN-based methods, indicating that more advanced techniques are required to fully exploit graph structure.
Comparing with reranker-based methods. gte-large achieves the second-best performance on both the ExplaGraphs and WebQSP datasets (F1 metric), outperforming other reranker approaches. However, it still falls significantly short of Align-GRAG. This demonstrates that simply performing node and triple reranking is insufficient to fully capture the graph structural information. Although G-RAG-RL also utilizes a GNN encoder to model graph structures, it relies on similarity with the query as a label, resulting in poor performance. This validates the effectiveness of our approach, which uses reasoning chains summarized by LLM as an optimization target. By aligning graph reasoning with LLM-derived reasoning chains, Align-GRAG achieves superior performance, showing that this alignment is crucial for graph-based reasoning tasks.
Comparing with GNN-based methods. G-Retriever ranks second on SceneGraphs, while GNP achieves the second-best Accuracy on WebQSP. However, both methods fall significantly short of Align-GRAG overall. This is because G-Retriever simply concatenates graph embeddings with LLM inputs using a projector, while GNP employs cross-modality pooling to fuse text and graph representations. Despite these efforts, their approaches remain limited. In contrast, Align-GRAG explicitly aligns the representation of graphs with language, effectively bridging the gap between these two spaces and achieving state-of-the-art results.
5.5. Ablation Study
Figure 3. Ablation study with different alignment strategy.
We perform an ablation study on WebQSP using three metrics to evaluate how the proposed modules contribute to performance improvement. As illustrated in Figure 3, we examine the following variants: (1) w/o Node Alignment: Removes the KL loss and knowledge pruning components. (2) w/o Representation Alignment: Excludes the contrastive loss and the graph embedding input to the LLM. (3) w/o Both: Removes both Node Alignment and Representation Alignment. (4) Random Alignment: Optimizes the aligner module using randomly generated labels. We observe that removing the Node Alignment module significantly lowers all evaluation metrics compared to the full model (Align-GRAG), indicating the effectiveness of node-level alignment optimization and node pruning. In contrast, while excluding the Representation Alignment module also leads to a performance decline, the impact is comparatively less significant. Performance further deteriorates when both modules are removed (w/o Both), highlighting the critical role of dual alignment between node and representation in filtering irrelevant knowledge and bridging the representation gap between the graph and language. The Random Alignment variant performs the worst, as random alignment fails to guide the model effectively. This emphasizes the necessity of meaningful alignment strategies and demonstrates that our LLM-summarized reasoning chain provides a highly useful label.
5.6. Generalization Analysis
Table 3. Generalization Analysis on different LLM and GNN backbones.
LLM Backbones Llama-2-7b-hf Llama-2-13b-hf Llama-3.1-8B
GNN Method ExplaGraphs WebQSP ExplaGraphs WebQSP ExplaGraphs WebQSP

### v3

@@ 3.2 Guided Dual Alignment
We propose an Aligner module that utilizes the extracted anchors to identify and align relevant nodes within the graph, effectively pruning redundant nodes and edges by filtering out unrelated information at the node level. Specifically, we employ a GNN (e.g., GraphTransformer Shi et al. (2020) or GAT Veličković et al. (2018)), to encode the structural information of the graph. The GNN produces node-level embeddings $\boldsymbol{n}_{g}$ based on the input graph:
$\boldsymbol{n}_{g}=\text{GNN}(\mathcal{G})\in\mathbb{R}^{|\mathcal{V}|\times d},$ (1)
@@
where $|\mathcal{V}|$ is the number of node and $d$ is feature dim. For the anchors, we employ SBERT to encode the textual description into a embedding $\boldsymbol{r}_{s}$ that captures its semantic meaning:
$\boldsymbol{r}_{s}=\text{SBERT}(t_{\text{anchor}})\in\mathbb{R}^{d_{s}}.$ (2)
@@
We concatenate the embedding of each node in the subgraph with the query embedding. The concatenated embeddings are then passed through an MLP module, which generates a predicted importance score for each node $\boldsymbol{p}_{\text{prediction}}$. These scores are transformed into probability distributions using the softmax function, which produces importance scores for both the prediction and the anchor.
$\boldsymbol{p}_{\text{prediction}}$ $=\text{Softmax}(\text{MLP}([\boldsymbol{n}_{g},\boldsymbol{q}^{\text{expand}}]))\in\mathbb{R}^{|\mathcal{V}|},$ (3)
$\boldsymbol{p}_{\text{anchor}}$ $=\text{Softmax}(\text{cos}(\boldsymbol{n}_{t},\boldsymbol{r}_{s}))\in\mathbb{R}^{|\mathcal{V}|},$ (4)
@@
where $[,]$ means concat operation, cos() means cosine similarity, $\boldsymbol{n}_{t}$ is text embedding of node, $\boldsymbol{q}^{\text{expand}}$ is the query embedding broadcasted across all nodes. To align the node distribution, we minimize the Kullback–Leibler (KL) divergence Kullback and Leibler (1951) between predicted probabilities $\boldsymbol{p}_{\text{prediction}}$ and the probabilities $\boldsymbol{p}_{\text{anchor}}$. The KL divergence loss for a subgraph is given by:
$\mathcal{L}_{\text{NA}}=\frac{1}{|\mathcal{V}|}\sum_{i=1}^{|\mathcal{V}|}\boldsymbol{p}_{\text{anchor}}(i)\log\frac{\boldsymbol{p}_{\text{anchor}}(i)}{\boldsymbol{p}_{\text{prediction}}(i)}.$ (5)
@@
Graph-level Alignment.
To bridge the representation gap between graph structures and textual descriptions, our Aligner module treats the text representation derived from rationale chains as the target label, aligning the graph and text embeddings to encourage semantic consistency. We apply a mean pooling operation across the node embeddings $\boldsymbol{n}_{g}$ to obtain a unified graph-level token $\boldsymbol{r}_{g}$:
$\boldsymbol{r}_{g}=\text{POOL}(\boldsymbol{n}_{g})=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\boldsymbol{n}_{g}(v)\in\mathbb{R}^{d}.$ (6)
@@
$\boldsymbol{\hat{r}}_{s}=\text{MLP}(\boldsymbol{r}_{s}),\quad\boldsymbol{\hat{r}}_{g}=\text{MLP}(\boldsymbol{r}_{g}).$ (7)
@@
The contrastive loss for graph-level alignment from $\boldsymbol{\hat{r}}_{g}$ to $\boldsymbol{\hat{r}}_{s}$ is defined as:
$\mathcal{L}_{GA(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s})}=-\frac{1}{N}\sum_{i=1}^{N}\left[\log\frac{\exp\left(\text{cos}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{cos}(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{j})/\tau\right)}\right],$ (8)
@@
where, $N$ is the batch size, $(\boldsymbol{\hat{r}}_{g}^{i},\boldsymbol{\hat{r}}_{s}^{i})$ is the $i$-th positive (graph-text) pair in the batch, $\tau$ is a temperature parameter to control the sharpness. Similarly, we can obtains the loss $\mathcal{L}_{GA(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})}$ from $\boldsymbol{\hat{r}}_{s}$ to $\boldsymbol{\hat{r}}_{g}$. The final representation alignment loss is obtained by:
$\mathcal{L}_{GA}=\frac{1}{2}\Big(\mathcal{L}_{GA(\boldsymbol{\hat{r}}_{s}\rightarrow\boldsymbol{\hat{r}}_{g})}+\mathcal{L}_{GA(\boldsymbol{\hat{r}}_{g}\rightarrow\boldsymbol{\hat{r}}_{s})}\Big).$ (9)
@@
To achieve an optimized Graph Aligner, we perform joint optimization of node and graph Alignment. The total loss for the Graph Aligner is defined as: $\mathcal{L}_{\text{Aligner}}=\mathcal{L}_{\text{GA}}+\mathcal{L}_{\text{NA}}$. The parameters of GNN encoder are jointly optimized using the loss $\mathcal{L}_{\text{Aligner}}$ with a specified training step, which reflects degree of alignment. We evaluate the impact of the alignment degree in Figure 3 and the effectiveness in Section 4.5.
3.2.1 Graph Token and Pruning
We apply the trained Graph Aligner to obtain an aligned graph representation token and a pruning subgraph for generation. Note that gold answers are only used to train the Aligner.
Encoding graph token.
Given the retrieved subgraph, we leverage the Aligner to produce node embeddings and a pooled graph token $\boldsymbol{r}_{g}$, which is optimized by $\mathcal{L}_{\text{Aligner}}$ to be consistent with the rationale-derived text space.
Structure-aware pruning.
We then use the node-level prediction scores to remove irrelevant evidence. Specifically, we select the top $n_{\text{seed}}$ nodes as seed nodes and expand them with their first-order neighbors to form the pruned subgraph $\mathcal{G}_{\text{Aligner}}=(\mathcal{V}_{\text{Aligner}},\mathcal{E}_{\text{Aligner}})$.
Finally, we concatenate the linearized text of $\mathcal{G}_{\text{Aligner}}$ with the query tokens $t_{q}$ and feed them into the LLM embedding layer to obtain token embeddings $\boldsymbol{r}_{t}$. We fuse $\boldsymbol{r}_{t}$ with $\boldsymbol{r}_{g}$ to condition answer generation:
@@ 3.2.1 Graph Token and Pruning
$p_{\theta}(Y\mid t_{q},\mathcal{G}_{\text{Aligner}})=\prod_{i=1}^{m}p_{\theta}\!\left(y_{i}\mid[\boldsymbol{r}_{g},\boldsymbol{r}_{t}],\,y_{<i}\right).$ (10)
@@
For efficiency, we fine-tune the generator with parameter-efficient methods such as LoRA Hu et al. (2021).
4 Experiments
4.1 Experimental Settings
@@
Table 1: Performance comparison using the same retrieval settings and generator for all methods. The table reports the mean and standard deviation results across three random seeds. For methods marked with $\dagger$’, we reproduce the results on WebQSP to report F1 and Accuracy. Results for methods marked with $\ddagger$’ are taken directly from He et al. (2024). The best results are highlighted in bold, and the second-best results are underlined. ‘Improvement $(\triangle)$’ represents the gain over second-best baseline. ‘*’ indicates statistically significant improvements (i.e., two-sided t-test with $p<0.05$) over second-best baseline. $\uparrow$: higher is better.
Datasets.
Following G-Retriever He et al. (2024), we conducted experiments on the GraphQA benchmark He et al. (2024), which includes ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). The dataset statistics are summarized in Appendix B.1.
Baselines.
We group baselines into four categories. (1) Inference-only methods leverage frozen LLMs by taking textualized graph and query as input, including Zero-shot, Zero-CoT Wei et al. (2022), CoT-BAG Wang et al. (2024), KAPING Baek et al. (2023), and ToG Sun et al. . (2) Raw Fine-tuning adapts LLM via parameter-efficient tuning, covering Prompt Tuning Lester et al. (2021) and LoRA Hu et al. (2021). (3) Reranker-based baselines rerank retrieved documents before feeding them into LLM, including GTE Li et al. (2023), BGE Xiao et al. (2024), G-RAG, and G-RAG-RL Dong et al. (2024). (4) GNN-based approaches integrate GNN encoders with LLM embeddings, including GraphToken Perozzi et al. (2024), GNP Tian et al. (2024), G-Retriever He et al. (2024), and GRAG Hu et al. (2025).
Implementation Details.
We use GraphTransformer Shi et al. (2020), GAT Veličković et al. (2018), and GCN Kipf and Welling (2016) as GNN encoders, and Llama-2-7b-hf, Llama-2-13b-hf and Llama-3.1-8b Touvron et al. (2023) as generators. For the retrieval process, we use the same retrieval results across all baselines to ensure fair comparisons. Both reranker-based and GNN-based methods apply LoRA for fine-tuning. All methods are compared using the same training hyperparameters where applicable. In our module, we explore two hyperparameters: alignment degree and the number of seed nodes ($n_{\text{seed}}$), with the analysis shown in Figure 8. To extract the anchor and rationale, we employ Llama-3.1-70B-Instruct Dubey et al. (2024). More details can be found in Appendix B.
LLM Backbones Llama-2-7b-hf Llama-2-13b-hf Llama-3.1-8B Qwen3-8B
GNN Method ExplaGraphs WebQSP ExplaGraphs WebQSP ExplaGraphs WebQSP ExplaGraphs WebQSP
$uparrow Accuracy\uparrow Hit@1\uparrow$ $uparrow Accuracy\uparrow Hit@1\uparrow$
GT GNP 0.8704 0.7391 0.8880 0.7696 0.8892 0.7557 0.8915 0.7602
G-Retriever 0.8705 0.7366 0.9115 0.7739 0.9014 0.7482 0.9088 0.7535
Align-GRAG 0.8992 0.7626 0.9241 0.7789 0.9321 0.7712 0.9365 0.7794
GAT GNP 0.9061 0.7291 0.8989 0.7676 0.9062 0.7521 0.9104 0.7588
G-Retriever 0.7960 0.7414 0.8953 0.7737 0.9111 0.7719 0.9150 0.7760
Align-GRAG 0.9151 0.7309 0.9151 0.7573 0.9214 0.7742 0.9282 0.7815
GCN GNP 0.7545 0.7298 0.8682 0.7564 0.8572 0.7495 0.8615 0.7530
G-Retriever 0.8592 0.7352 0.9007 0.7521 0.8512 0.7527 0.8588 0.7592
Align-GRAG 0.8574 0.7377 0.9152 0.7592 0.8596 0.7573 0.8695 0.7645
Table 2: Generalization Analysis on different LLM and GNN backbones.
4.2 Main Results
We conduct extensive experiments against 18 baselines, and Table 1 shows that Align-GRAG consistently improves all metrics across three datasets, with a particularly large Accuracy gain on WebQSP over the second-best method due to its alignment-based graph pruning and optimization. Compared with inference-only methods, which rely purely on LLM reasoning (e.g., Zero-shot) or heuristic KG search (e.g., ToG), Align-GRAG benefits from task-aware structural optimization and achieves clearly stronger results. Compared with raw fine-tuning methods such as Prompt Tuning and LoRA, Align-GRAG further demonstrates that parameter-efficient adaptation alone is insufficient to fully exploit graph structure. Compared with reranker-based methods (e.g., gte-large and related variants), which mainly rerank nodes/triples and thus miss richer structural dependencies, Align-GRAG achieves superior performance by aligning graph reasoning with LLM-derived reasoning chains as the optimization target. Compared with GNN-based methods (e.g., G-Retriever and GNP), which inject graph embeddings via projection or pooling but lack explicit graph–language alignment, Align-GRAG more effectively bridges the two representation spaces and yields state-of-the-art performance.
4.3 Generalization Analysis
In this section, we evaluate whether Align-GRAG generalizes across both GNN architectures and LLM scales on ExplaGraphs (Accuracy) and WebQSP (Hit@1). As summarized in Table 2, across three GNN backbones (GT, GAT, GCN) and multiple LLM families/sizes, Align-GRAG outperforms GNP and G-Retriever in the vast majority of configurations, indicating strong robustness to different encoder–reasoner pairings. The largest gains are observed with the GraphTransformer backbone, suggesting that its ability to capture long-range dependencies synergizes with Align-GRAG ’s alignment objectives to improve retrieval and reasoning, while consistent improvements with GAT further show that alignment complements attention-based aggregation. Moreover, scaling up the LLM generally boosts results—most notably with the weaker GCN encoder—implying that larger LLMs can better leverage aligned graph signals to compensate for limited graph expressivity. The consistent gains on both datasets confirm that Align-GRAG transfers well across different settings.
4.4 Ablation Study
We conduct an ablation study on WebQSP with three metrics to assess the contribution of each module. As shown in Table 3, we consider four variants: (1) w/o Node Alignment (removing the KL loss and pruning), (2) w/o Graph Alignment (removing the contrastive loss and graph token), (3) w/o Both, and (4) Random Alignment (training the aligner with random labels). Removing Node Alignment causes the largest degradation across all metrics, confirming the importance of node-level alignment and pruning. Disabling Graph Alignment also reduces performance, and removing both modules further amplifies the drop, underscoring the necessity of dual alignment for suppressing irrelevant knowledge and mitigating the graph language representation gap. Finally, Random Alignment performs worst, indicating that effective alignment requires meaningful supervision; our LLM-extracted anchor and rationale provide such a useful signal.
Method WebQSP